a rectangle is twice as long as it is wide its area is 450 sq centimeters,what are the dimentions?
Answer by mananth(16946)

a rectangle is twice as long as it is wide its area is 450 sq centimeters,what are the dimentions?
Let the width be x
The length will be 2x
Area of a rectangle is given by L * W
2x*x = 450 sq. cms
2x^2= 450
x^2=225
x+ sqrt 225
The width is 15 cms
The length is 30 cms